# Controlling photon-magnon coupling using another magnon

### v1

@@ Abstract
We investigate photon-mediated magnon–magnon coupling between spatially separated Yttrium Iron Garnet (YIG) and permalloy (NiFe) thin films placed on a planar hexagonal ring resonator. We observe clear signatures of magnon-magnon interaction in the absence of direct dipolar interaction between the magnetic films. Notably, the coupling strength between the hexagonal ring resonator and the permalloy film increases with the thickness of the YIG film, despite a fixed permalloy film thickness, suggesting the presence of an indirect interaction channel mediated by resonator photons. To support these findings, we present a theoretical model that accurately reproduces the observed transmission spectra and reveals a nontrivial interdependence between the individual coupling strengths of YIG and permalloy to the resonator. These results underscore the importance of indirect interactions and potential crosstalk pathways in designing hybrid magnonic systems and scalable quantum architectures, while demonstrating the feasibility of cost-effective, planar configurations for experimental implementation. These insights are valuable for advancing low-loss, coherent information transfer in hybrid quantum devices.
@@
Unresolved theoretical and experimental challenges persist, particularly in understanding mediated coupling, the dependence of interaction strength on spin density and spatial configuration, and the role of decoherence in nonlocal systems. Notably, Hyde et al. [19] observed that specific indirectly coupled modes can surpass their directly coupled counterparts in transmission amplitude, highlighting the need to reassess the conventional assumptions surrounding circuit isolation and crosstalk. As previous work suggests that cavity photons can act as effective mediators for long-distance coupling in hybrid architectures [20], we investigate this phenomenon using a hexagonal ring resonator as the coupling platform. It is planar and symmetric: well-suited for probing such mediated interactions and exploring their impact on the coherent dynamics of spatially separated magnetic elements.
@@
In designing hybrid systems, it is essential to account for the various forms of hybridization and their collective influence on the system’s dynamics. We investigate magnonic interactions in a configuration specifically chosen to minimize direct dipolar coupling. This allows us to isolate and analyze interaction pathways mediated indirectly through a standard bosonic mode, i.e., photons. The observed hybrid system features the resonator parked near a microstrip line and integrated with the two magnetic films—YIG and permalloy—mounted on opposite arms of the resonator within a planar geometry, as illustrated in Fig. 1(a). When a microwave current propagates along the -axis through the microstrip line, it produces a transverse microwave magnetic field that excites the hexagonal resonator (depicted as the silver strip on the substrate). It functions as a parallel-LC circuit exhibiting quasi-static resonance. This resonant field, in turn, interacts with the magnetic YIG (black film in Fig. 1(a)) and permalloy (dirty green film) films placed on the resonator, inducing their respective magnon modes and enabling the probing of respective photon–magnon coupling dynamics. The high-frequency electromagnetic solver CST Studio Suite was employed to simulate individual hybrid configurations: resonator–permalloy, resonator–YIG, and the combined hybrid system resonator–YIG–permalloy. The simulations employed standard material parameters for the magnetic films and are listed in Table 1. The dimensions of the resonator, the microstrip line, the YIG and permalloy films are indicated in Table 2 (cf. Verma et al.[32]).
@@
To investigate photon-mediated magnon–magnon coupling, we obtained as a function of microwave frequency () under a static external magnetic field () applied along the -axis at room temperature. Figures 2a, 2b, and 2c show the power spectra on the – plane for the hybrid systems resonator–permalloy, resonator–YIG, and resonator–YIG–permalloy, respectively. In this setup, the photon mode associated with the resonator behaves as a fixed-frequency resonance determined by the geometry of the LC structure. In contrast, the magnon modes in YIG and permalloy follow the Kittel relation[33], resulting in a resonance that shifts with the applied magnetic field. The strength of photon-magnon coupling is inferred from the degree of avoided level crossing–an established indicator of coherent energy exchange–and is a particularly relevant parameter in applications such as quantum sensing and quantum memory buses[15].
@@
As shown in Fig. 2a, the permalloy–resonator system exhibits only weak coupling, which is expected due to the metallic nature of permalloy that introduces higher damping and noise. In our model, we observe a coupling constant of corresponding to this interaction. In contrast, Fig. 2b demonstrates that the YIG–resonator system exhibits strong coupling (), attributable to the low damping and high spin-wave coherence of the magnetic material involved. Interestingly, in the combined system (Fig. 2c), the permalloy mode displays a markedly enhanced coupling signature compared to its isolated behavior, while the YIG coupling appears slightly diminished. We note that fitted values are now (of permalloy) and (of YIG), a simultaneous enhancement of permalloy-resonator coupling (seen at crossing marked P1 in Fig. 2c) and reduction of YIG-resonator coupling (crossing P2). This suggests a redistribution of interaction strength facilitated by the resonator’s photon mode, emphasizing the role of photon mediation in enabling effective magnon–magnon interactions between spatially separated magnetic elements.
@@
Fig. 1(b) attempts to visually illuminate the nature of this model. Three oscillators named 1, 2, and , each with its own internal damping and a coupling to travelling photons, interact with one another. While oscillators 1 and 2 interact with oscillator directly, they are not allowed to interact with one another.
@@
and captures the crosstalk between the resonator modes, leading to the observed normal anticrossing in Fig. 2. The normal anticrossing behaviour is due to real coupling constants in the Hamiltonian [34, 35].
@@
To investigate the presence of indirect magnon–magnon coupling mediated by cavity photons, we performed a series of simulations by varying the thickness () of the YIG film (from to m) while keeping the properties of the permalloy film and the hexagonal microwave resonator unchanged. Increasing the YIG thickness enhances its spin density, thereby modifying its magnonic response.
@@
As shown in the top row of Fig. 3, an increase in the mode splitting between the upper and lower branches (for both regions P1 and P2) of the magnon–photon hybrid modes is observed with thicker YIG films. Notably, this occurs when the permalloy film remains unaltered, indicating that the enhanced coupling is not due to changes in itself. We make a more direct observation of this in the plots in Fig. 4, where we compare the coupling of the resonator-YIG in the absence of the permalloy film in Fig. 4(a) and in its presence in Fig. 4(b) (blue crosses). Both linear.
@@
Evidently, while the slope of the resonator-YIG coupling remains unaffected, the intercept drops slightly in the presence of another magnon. Within the region of interest, the presence of the permalloy film causes a constant decrease in resonator-YIG coupling. Combined with the non-zero slope of the unvarying resonator-permalloy couple, the behavior provides strong evidence for photon-mediated interaction between the YIG and permalloy magnon modes, a hallmark of indirect magnon–magnon coupling through the shared cavity photon field.
@@
To validate these observations, we used a theoretical model to compute the transmission spectrum () under identical conditions. The bottom row of Fig. 3 presents the calculated spectra, which closely match the corresponding observed results in the top row. The evolution of the spectra with increasing YIG thickness—from left to right—demonstrates consistent hybrid dynamics, affirming the model’s accuracy. Predictably, as YIG becomes thicker, its coupling to the resonator increases due to the higher spin density. However, a notable and somewhat counterintuitive result is the simultaneous increase in the coupling strength of the permalloy mode, despite its parameters remaining fixed. This reveals that the YIG film exerts an indirect influence on the permalloy–resonator coupling, mediated by their mutual interaction with the cavity mode. We can ascribe to the resonator-permalloy coupling a dependence on the resonator-YIG coupling, , which we observe in this work to be linear (Fig. 4(c)).
Parameter | Value |
YIG Sat. Mag. () | 1750 G |
YIG Gyromag. Ratio () | |
permalloy Sat. Mag. () | 10900 G |
permalloy Gyromag. Ratio () |
Parameter | Value (mm) |
Thin film side length (permalloy & YIG) | 3 |
permalloy thickness | 0.02 |
Resonator outer radius | 4 |
Resonator inner radius | 3.4 |
MSL width | 0.57 |
Substrate length | 30 |
Substrate width | 20 |
Substrate thickness | 0.64 |
Copper thickness (MSL and ground plane) | 0.035 |
In conclusion, this work bridges the gap between fundamental magnonics and practical device engineering. Elucidating photon-mediated coupling in accessible macroscopic systems provides a roadmap for low-cost classical applications and future quantum technologies requiring precise control over indirect spin-photon interactions.
Acknowledgements.
@@
The work was supported by the Council of Science and Technology, Uttar Pradesh (CSTUP) (Project Id: 2470, CST, U.P. sanction No: CST/D-1520 and Project Id: 4482, CST, U.P. sanction No: CST/D-7/8). B. Bhoi acknowledges support from the Science and Engineering Research Board (SERB) India - SRG/2023/001355. S. Verma acknowledges the Ministry of Education, Government of India, for the Prime Minister’s Research Fellowship (PMRF ID-1102628).Statements and Declarations

### v3

# Dynamic control of photon-magnon interactions via secondary magnon excitation

@@ Abstract
Photon-mediated magnon-magnon coupling between spatially separated Yttrium Iron Garnet (YIG) and permalloy (NiFe) thin films on a planar hexagonal ring resonator shows clear signatures of magnon-magnon interaction are observed without direct dipolar interaction between the magnetic films. The coupling strength between the hexagonal ring resonator and the permalloy film increases with the thickness of the YIG film, despite a fixed permalloy film thickness. This suggests the presence of an indirect interaction channel mediated by resonator photons. A theoretical model is presented that accurately reproduces the observed transmission spectra and reveals a nontrivial interdependence between the individual coupling strengths of YIG and permalloy to the resonator. These results highlight the importance of indirect interactions and potential crosstalk pathways in designing hybrid magnonic systems and scalable quantum architectures, while demonstrating the feasibility of cost-effective, planar configurations for experimental implementation. These insights are valuable for advancing low-loss, coherent information transfer in hybrid quantum devices.
keywords:
Photon, magnon, indirect coupling, hybrid quantum device, coherentFizaan Khan111fizaan.khan.phy21@iitbhu.ac.in Sachin Verma Biswanath Bhoi Rajeev Singh222rajeevs.phy@iitbhu.ac.in
Nano-Magnetism and Quantum Technologies Laboratory
Indian Institute of Technology (BHU), Varanasi
221005 - India
1 Introduction
@@
Unresolved theoretical and experimental challenges persist, particularly in understanding mediated coupling, the dependence of interaction strength on magnetization and spatial configuration, and the role of decoherence in nonlocal systems. Notably, Hyde et al. [19] observed that specific indirectly coupled modes can surpass their directly coupled counterparts in transmission amplitude, highlighting the need to reassess the conventional assumptions surrounding circuit isolation and crosstalk. As previous work suggests that cavity photons can act as effective mediators for long-distance coupling in hybrid architectures [20], we investigate this phenomenon using a hexagonal ring resonator as the coupling platform. It is planar and symmetric: well-suited for probing such mediated interactions and exploring their impact on the coherent dynamics of spatially separated magnetic elements.
@@
2 Design and Numerical Modeling of the Hybrid Photon-Magnon System
@@
In designing hybrid systems, it is essential to account for the various forms of hybridization and their collective influence on the system’s dynamics. We investigate magnonic interactions in a configuration specifically chosen to minimize direct dipolar coupling. This allows us to isolate and analyze interaction pathways mediated indirectly through a standard bosonic mode, i.e., photons. The observed hybrid system features the resonator parked near a microstrip line that is 0.57 mm wide and 30 mm long. It integrates with the two 3 mm 3 mm magnetic films—YIG and permalloy—mounted on opposite arms of the resonator within a planar geometry, as illustrated in Figure 1(a). When a microwave current propagates along the -axis (see axes in Figure 1(a)) through the microstrip line (labelled ‘MSL’ in Figure 1(a)), it produces a transverse microwave magnetic field that excites the hexagonal resonator (labelled ‘HRR’). It functions as a parallel-LC circuit exhibiting quasi-static resonance. This resonant field, in turn, interacts with the magnetic YIG (black film in Figure 1(a)) and permalloy (of thickness 20 m, dirty green film in the Figure) films placed on the resonator, inducing their respective magnon modes and enabling the probing of respective photon–magnon coupling dynamics. The high-frequency electromagnetic solver CST Studio Suite was employed to simulate individual hybrid configurations: resonator–permalloy, resonator–YIG, and the combined hybrid system resonator–YIG–permalloy.
@@
To investigate photon-mediated magnon–magnon coupling, we obtained as a function of microwave frequency () under a static external magnetic field () applied along the -axis at room temperature. Figures 2a, 2b, and 2c show the power spectra on the – plane for the hybrid resonator–permalloy, resonator–YIG, and resonator–YIG–permalloy, respectively. In this setup, the photon mode associated with the resonator behaves as a fixed-frequency resonance determined by the geometry of the LC structure. In contrast, the magnon modes in YIG and permalloy follow the Kittel relation[32], resulting in a resonance that shifts with the applied magnetic field. The strength of photon-magnon coupling is inferred from the degree of avoided level crossing–an established indicator of coherent energy exchange–and is a particularly relevant parameter in applications such as quantum sensing and quantum memory buses[15].
@@
As shown in Figure 2a, the permalloy–resonator system exhibits only weak coupling, which is expected due to the metallic nature of permalloy that introduces higher damping and noise. In our model, we observe a coupling constant of corresponding to this interaction. In contrast, Figure 2b demonstrates that the YIG–resonator system exhibits strong coupling (), attributable to the low damping and high spin-wave coherence of the magnetic material involved. Outside the influence of the magnonic film’s resonance behavior, the photonic resonator returns to its natural frequency, observed in the rise of the horizontal line in Figure 2b.
Interestingly, in the combined system (Figure 2c), the permalloy mode displays a markedly enhanced coupling signature compared to its isolated behavior, while the YIG coupling appears slightly diminished. We note that fitted values are now (of permalloy) and (of YIG), a simultaneous enhancement of permalloy-resonator coupling (seen at crossing marked P1 in Figure 2c) and reduction of YIG-resonator coupling (crossing P2). This suggests a redistribution of interaction strength facilitated by the resonator’s photon mode, emphasizing the role of photon mediation in enabling effective magnon–magnon interactions between spatially separated magnetic elements.
3 Theoretical Formalism for Photon Mediated Magnon-Magnon Coupling
@@
Figure 1(b) attempts to visually illuminate the nature of this model. Three oscillators named 1, 2, and , each with its own internal damping and a coupling to travelling photons, interact with one another. While oscillators 1 and 2 interact with oscillator directly, they are not allowed to interact with one another.
@@
and captures the crosstalk between the resonator modes, leading to the observed normal anticrossing in Figure 2. The normal anticrossing behaviour is due to real coupling constants in the Hamiltonian [33, 34].
@@ 4 Discussion
To investigate the presence of indirect magnon–magnon coupling mediated by cavity photons, we performed a series of simulations by varying the thickness () of the YIG film (from to m) while keeping the properties of the permalloy film and the hexagonal microwave resonator unchanged. Increasing the YIG thickness enhances its magnetization, thereby modifying its magnonic response.
@@
As shown in the top row of Figure 3, an increase in the mode splitting between the upper and lower branches (for both regions P1 and P2) of the magnon–photon hybrid modes is observed with thicker YIG films. Notably, this occurs when the permalloy film remains unaltered, indicating that the enhanced coupling is not due to changes in itself. We make a more direct observation of this in the plots in Figure 4, where we compare the coupling of the resonator-YIG in the absence of the permalloy film in Figure 4a and in its presence in Figure 4b (blue crosses). Both linear.
@@
Evidently, while the slope of the resonator-YIG coupling (i.e., vs ) remains unchanged from Figure 4b to Figure 4a, the intercept does drop slightly. The presence, then, of the permalloy film causes a decrease in resonator-YIG coupling independent of the thickness of the YIG film. Combined with the non-zero slope of the unvarying resonator-permalloy couple, the behavior provides strong evidence for photon-mediated interaction between the YIG and permalloy magnon modes, a hallmark of indirect magnon–magnon coupling through the shared cavity photon field.
@@
To validate these observations, we used a theoretical model to compute the transmission spectrum () under identical conditions. The bottom row of Figure 3 presents the calculated spectra, which closely match the corresponding observed results in the top row. The evolution of the spectra with increasing YIG thickness—from left to right—demonstrates consistent hybrid dynamics, affirming the model’s accuracy. Predictably, as YIG becomes thicker, its coupling to the resonator increases due to the higher magnetization. However, a notable and somewhat counterintuitive result is the simultaneous increase in the coupling strength of the permalloy mode, despite its parameters remaining fixed. This reveals that the YIG film exerts an indirect influence on the permalloy–resonator coupling, mediated by their mutual interaction with the cavity mode. We can ascribe to the resonator-permalloy coupling a dependence on the resonator-YIG coupling, , which we observe in this work to be linear (Figure 4c).
5 Conclusions
In this study, we have demonstrated and quantified the photon-mediated coupling between two spatially separated magnonic films, YIG and permalloy, integrated onto a planar hexagonal ring resonator. Our experimental results reveal that even without direct dipolar interaction, a clear magnon-magnon coupling emerges, evidenced by the anticrossing in the transmission spectra. A key finding is the non-trivial enhancement of the permalloy-resonator coupling strength as the thickness of the YIG film increases, while the permalloy film itself remains unchanged. This observation strongly indicates that the resonator’s photonic mode acts as a quantum bus, facilitating an indirect interaction pathway between the two magnonic systems.
Furthermore, the theoretical model developed using the input-output formalism accurately reproduces the experimental transmission spectra across various YIG thicknesses. The model confirms a linear relationship between the coupling strengths of the YIG and permalloy films to the resonator, quantitatively capturing the indirect influence one magnon has on the other. These findings underscore the critical role of mediated interactions in hybrid magnonic systems, offering a new mechanism for controlling coupling strengths in scalable, on-chip architectures and paving the way for advanced quantum and microwave technologies.
Acknowledgements
@@
The work was supported by the Council of Science and Technology, Uttar Pradesh (CSTUP) (Project Id: 2470, CST, U.P. sanction No: CST/D-1520 and Project Id: 4482, CST, U.P. sanction No: CST/D-7/8). B. Bhoi acknowledges support from the Science and Engineering Research Board (SERB) India - SRG/2023/001355. S. Verma acknowledges the Ministry of Education, Government of India, for the Prime Minister’s Research Fellowship (PMRF ID-1102628).
Statements and Declarations
@@
Appendix
The following is a detailed derivation of the transmission parameter , using the input-output formalism[28, 36].
The Hamiltonian of the system was asserted in Section 3 to exist in three parts: (1), (2), and (3). Plugging these into and making the rotating wave approximation we obtain
The dynamics of the system can be described by the Heisenberg-Langevin equations of motion, which are derived from the Hamiltonian and the fundamental commutation relations for bosonic operators that have the commutation relations and .
The solutions to the Heisenberg equations of motion for and for some are the following:
Whereas, when the solution is sought until some :
We pull the out of the integral under the assumption that the transfer of energy between the operators and is a Markov process. For :
We will then define and as
We can now define the decay rate . This allows us to further simplify the equations of motion by plugging the equations of the input or output modes (11) into (10) to yield
where is defined as
Taking the Fourier transform of these expanded equations gives
Finally, by defining the parameter as and substituting the derived expressions, we arrive at the final result:
which matches Equation 4 precisely.